OpenStudy (anonymous):

Which of the following is an equation of a line passing through the points (-1,2) and (-3,-2)? A) y=1/2x+4 B) y/2=x+4 C) y=2(x+2) D) y-x/2=8

OpenStudy (***[isuru]***):

can u find the slope of the line using (-1,2) and (-3,-2) ?

OpenStudy (anonymous):

y=mx+b form?

OpenStudy (***[isuru]***):

\[slope = \frac{ y_{2} - y_{1} }{ x_{2} - x_{1} }\]

OpenStudy (anonymous):

that;s the formula, yes, but what is the slope?

OpenStudy (***[isuru]***):

yep ! if u want it that way u can find the slope and go to slope intercept form or y= mx + b or u could do it using point slope form.. but before that u must find the slope

OpenStudy (anonymous):

(-2)-2/(-3)--1

OpenStudy (gorv):

@Decayer_ FOOLOW @***[ISURU]***

OpenStudy (anonymous):

Is it B?

OpenStudy (anonymous):

answer i c

OpenStudy (anonymous):

y = 2x +4

OpenStudy (***[isuru]***):

\[( x_{1} \ \ \ \ , \ \ \ \ \ \ y_{1} ) \ \ \ ( x_{2} \ \ \ \ \ , \ \ \ \ y_{2})\] ( -1 , 2) ( -3 , -2) now plug them in to \[slope = \frac{ y_{2} - y_{1} }{ x_{2} - x_{1} }\]

OpenStudy (anonymous):

Or, from your answers it would be y= 2(x+2)

OpenStudy (anonymous):

and as suru said the formulae is for slope which is not compolsory to know

OpenStudy (***[isuru]***):

can u do that ?

OpenStudy (***[isuru]***):

Dageek did u found the slope ?

OpenStudy (anonymous):

C?